# Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Multi-modal Text Recognition

### v3

@@ 5 Discussion
Limitations. Despite promsing results, we identified three cases where GFD is most susceptible to failure, which is an exciting future direction. We categorize them into ASR errors and LLM errors.
@@
ASR proposed errors: When ASR proposes a candidate with high probability that deviates from the phonetic constraint, the LLM may falsely pickup the sequence, inducing an error. There are two main types - semantically activated tokens and time delayed activated tokens. Semantically activated tokens are tokens that are encoded in similar output embedding space due to semantic similarity. We found that these errors are much more common in Chinese and identified some pairs from a top-k analysis of a decoding sequence, e.g., {CJK}UTF8bsmi’剛’ and’才’, ’做(什麼)’ and ’幹(什麼)’ . Time delay activated tokens are proposed tokens that are targets at a later step, where selecting them effectively skips some intermediate tokens. We observe that these tokens are much more likely to be present at the start of the sequence.
@@ 6 Conclusion
The Generative Fusion Decoding (GFD) framework represents a leap forward in fusing auto-regressive text models from heterogeneous token spaces. Our experiments confirm GFD’s effectiveness, particularly in ASR tasks. For tonal languages like Chinese, it achieves a mixed error rate of 8.83% on the NTUML2021 dataset, surpassing previous benchmarks. In context-aware recognition, GFD demonstrated its prowess by a 17.7% Word Error Rate Reduction (WERR) on the ATCO2 dataset when utilizing prompts to both ASR and LLM . For OCR tasks, GFD’s fusion with LLMs resulted in a notable 38.06% improvement in exact matches on the NAF-Long dataset, underscoring the framework’s ability to enhance recognition of long text sequences. These experimental outcomes highlight GFD’s robustness and adaptability, showcasing its generality to leverage the extensive training and contextual understanding of LLMs to improve text recognition across various domains and challenges. We hope that our framework will motivate further research into more fusion techniques that better leverages the existing capabilties of the pre-trained models.

### v4

@@ 5.3 Error Analysis
Despite promising results, we identified three cases in which GFD is most susceptible to failure, representing exciting directions for future work. We categorize these failure cases into ASR errors and LLM errors.
@@
ASR proposed errors: When ASR proposes a candidate with high probability that deviates from the phonetic constraint, the LLM may falsely pickup the sequence, inducing an error. There are two main types - semantically activated tokens and time delayed activated tokens. Semantically activated tokens are tokens that are encoded in similar output embedding space due to semantic similarity. We found that these errors are much more common in Mandarin. Time delay activated tokens are proposed tokens that are targets at a later step, where selecting them effectively skips some intermediate tokens. We observe that these tokens are much more likely to be present at the start of the sequence.
@@ 6 Conclusion
In summary, we propose Generative Fusion Decoding (GFD), a simple yet effective framework for integrating large language models into ASR and OCR systems through byte-level shallow fusion. Our theoretical derivations provide a foundation for this integration, while empirical results demonstrate consistent gains across diverse conditions—including noisy audio, long-form inputs, and instruction-following tasks. GFD also compares favorably with prior fusion methods, highlighting its potential as a general-purpose solution. These results establish a foundation for future exploration of fusion methods that further exploit the strengths of pre-trained language models.
Limitations
The effectiveness of GFD is hindered when LLM selects an ASR token candidate that deviate from the correct phonetic content, leading to hallucinations. We provide in our analysis general categories of these errors, for practitioners to be aware of such a risk. We also advise users to carefully select the LLM, as the LLM itself may have limitations in its understanding or biases present in its training data. If the LLM misinterprets context or generates incorrect predictions, these errors can propagate through the GFD framework, affecting the overall performance.
Appendix A Appendix
A.1 Recursive Calculation of the GFD formula
We have shown the efficient calculation of $\mathbb{P}_{m,approx}(\{B_{l}\},\mathcal{Z}^{(m)})$ with equation (7). We now show that the incremental calculation of $\mathbb{P}_{m,approx}(\{B_{l}\},\mathcal{Z}^{(m)})$ from $\mathbb{P}_{m,approx}(\{B_{l-1}\},\mathcal{Z}^{(m)})$ is $O(1)$ in terms of the costly decoder forward.
Since the model forwarding is only dependent on $\hat{T}_{s}$, and not on alternative tokens, we first derive how the main sequence differs between $B_{l}$ and $B_{l-1}$. We denote the main sequence of $B_{l}$ as $\hat{T}_{s}$, and let the main sequence of $B_{l-1}$ as $\hat{T\textquoteright}_{s}$. In most scenarios, $\hat{T}_{s}$ is one of the two:
  • $\hat{T^{\prime}}_{s}$ plus one additional byte token.
  • The additional byte token merges with previous tokens in $\hat{T^{\prime}}_{s}$, a new token appends a truncated $\hat{T^{\prime}}_{s}$
In either of the cases, there will only be one additional token on the main path. Calculating alternative tokens only requires indexed selection through operations like "mask-select", which is inexpensive compared to the model forward operation.  [^3] The total FLOPs for the byte algorithm remain under $10^{5}$, which is negligible compared to model forwarding $(>10^{10})$ Therefore, with proper kv-caching on results of $B_{l-1}$, we can efficiently calculate $B_{l}$ to realize GFD.
A.2 Experimental Details on selecting $r$ in Equation 8
All GFD fused models are run on a single A6000 GPU. For the parameter of $r$ in Equation 8, we conduct grid search of among $[0.1,0.2,0.3,0.4]$ on noisy-librispeech, and selected $r=0.2$. We keep $r=0.2$ across all our experiments; while setting the number of beams equal to 5 or 10 for ASR and OCR experiments, respectively.
A.3 Experimental Details on selecting $k$ in Equation 8
To maintain reasonable time complexity, we aim to limit the number of rescoring samples to match the number of beams, i.e., num_beams. Assume num_beams=5. During the expansion phase of beam search, when $k=0$, the text recognition modality generates up to $5\times 5=25$ candidates, which is excessive. By strategically selecting $k$ such that the resulting sequence consists only of tokens from sequences prior to the expansion phase, the number of candidates will naturally be capped at the beam size. Thus, the optimal value of $k$ corresponds to the length of the last token proposed by the text recognition modality, varying across different beam hypotheses. Choosing a larger $k$ results in a loss of information, which is suboptimal.
A.4 Prompting Details
Here we list the prompting details of benchmarking.
Librispeech and noisy-librispeech
ASR Prompt: (None)
LLM Prompt: {mdframed} The following is a transcription of a spoken sentence:
Medical
ASR Prompt: (None)
LLM Prompt: {mdframed} The following is a transcription of a spoken sentence:
Fleurs-HK
ASR Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription: LLM Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription:
ML Lecture
ASR Prompt: {mdframed} (In Chinese) Traditional Chinese LLM Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription, there exists code-switching, and some of the vocabulary is in English.
Formosa
ASR Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription LLM Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription:
ATCO2
ASR Prompt: {mdframed} Alfa Bravo Charlie Delta Echo Foxtrot Golf Hotel India Juliett Kilo Lima Mike November Oscar Papa Quebec Romeo Sierra Tango Uniform Victor Whiskey Xray Yankee Zulu One Two Three Four Five Six Seven Eight Nine Zero
Dayton radio, November One Two Three Four Five on one two two point two, over Springfield V-O-R, over.
New York Radio, Mooney Three One One Echo. Columbia Ground, Cessna Three One Six Zero Foxtrot, south ramp, I-F-R Memphis. LLM Prompt: {mdframed} Section 2. Radio Communications Phraseology and Techniques
1. General
…(4000 words on call signs and regulations)…
Generative Error Correction
We follow Task-Activating Prompting method in Chen et al. (2024a) to create the prompt for Generative Error Correction. {mdframed} User: Do you know Automatic Speech Recognition?
Assistant: Yes, I do! …
User: Do you know language model restoring…
Assistant: Language model restoring is …
User: Can you generate an example with 5-best list?
Assistant: Sure! …
User: Please do the same thing on the following n-best list…